# Diffusion Model in Latent Space for Medical Image Segmentation Task

### v2

@@ Abstract
Medical image segmentation is essential for supporting clinical diagnosis and treatment planning. With growing imaging demands, AI models are increasingly used to improve the accuracy and efficiency of this task. However, traditional approaches often produce a single segmentation mask per input image, limiting the ability to capture uncertainty in the segmentation process. Recent advancements in generative models for image generation present new opportunities to address these limitations by enabling the generation of multiple segmentation masks for each input image. Nevertheless, the computational complexity and performance of these models remain challenging. In this work, we propose MedSegLatDiff, a novel model for medical image segmentation that integrates a diffusion-based model (DM) with a variational autoencoder (VAE) to decouple perceptual data compression and generation processes. The VAE compresses input data into a low-dimensional latent space, reducing noise and accelerating learning, while the DM performs segmentation more efficiently within this latent representation. Furthermore, we replace the traditional Mean Squared Error (MSE) loss with a weighted Cross-Entropy (WCE) loss in the perceptual mask compression module to enhance the reconstruction of segmentation masks during the encode–decode process, particularly for tiny nodules. This design emulates the collaborative segmentation approach of multiple doctors, offering a more robust and reliable performance compared to previous methods that replicate the segmentation patterns of individual practitioners. We evaluate our model on three datasets: ISIC-2018, CVC-Clinic, and LIDC-IDRI, demonstrating competitive performance while simultaneously providing the ability to generate confidence maps for deeper analysis by experts. Our results show that MedSegLatDiff achieves prominent segmentation performance while offering enhanced interpretability and consistency for clinical applications.
@@ I Introduction
Medical image segmentation involves partitioning a medical image into meaningful regions, playing a crucial role in diagnosis, surgical planning, and image-guided procedures. Accurate segmentation helps clinicians better interpret images, facilitate comparisons, and monitor changes over time. However, manually labeling segmentation masks is time-consuming and labor-intensive, motivating the development of automated methods. While classical techniques such as binary thresholding and k-means clustering have shown limited effectiveness, recent advances in artificial intelligence have enabled the widespread use of machine learning and deep learning models, achieving remarkable results. Typically, these AI models are trained on pre-labeled datasets and can automatically segment new images. Most existing approaches adopt a one-to-one strategy, producing a single segmented output per input image [2, 3, 23, 31]. This means generating only one mask per input, effectively replicating the behavior of a single doctor. However, this approach has inherent limitations when applied to complex and ambiguous data, particularly in medical contexts.
@@
Recently, the development of generative models, particularly diffusion-based models (DM) [5], has led to substantial advances in image generation and synthesis, often producing results superior to earlier approaches such as Generative Adversarial Networks (GAN) [19] and Variational AutoEncoder (VAE) [14], which remain among the most widely used methods in recent years. Beyond their success in general image generation, these models open up new possibilities for medical image segmentation through a one-to-many paradigm. Unlike traditional one-to-one approaches, diffusion-based models can generate multiple segmentation masks for a single input, capturing the model’s uncertainty in ambiguous medical data. Several recent studies have adopted this strategy and demonstrated superior performance and robustness compared to conventional methods [7, 8, 9, 26]. These results indicate that diffusion-based approaches not only enhance segmentation accuracy but also provide richer information to support clinical decision-making, particularly in challenging cases involving ambiguous or subtle anatomical structures. Moreover, we find that the model’s performance can still be further improved in terms of both speed and accuracy.
@@
In this paper, we propose MedSegLatDiff, a medical image segmentation model that follows a novel one-to-many approach based on DM. Our model incorporates a VAE and consists of two stages. The first performs an encoder–decoder operation between input data and latent representation, while the second stage carries out diffusion processes in the latent space to generate a set of segmentation masks for each input image. By imposing conditioning on the random generation process of the diffusion model in the latent space, the model can generate a set of segmentation masks whose variability captures the uncertainty typically exhibited by a group of doctors. Therefore, our model can simulate the segmentation process of a group of doctors and is well-suited for medical image segmentation tasks. At the same time, the synthesized results from the generated set of segmented images of the model can compete with previous models and confidence map assists doctors in making more in-depth diagnoses. To the best of our knowledge, the main contributions of this paper are as follows:
@@
  • Training and incorporating a VAE model to develop a conditional diffusion-based model in latent space, which reduces noise, compresses image size, and improves both training and inference efficiency.
  • Utilizing a weighted Cross-Entropy (CE) loss instead of Mean Squared Error (MSE) for tiny segmentation masks in the VAE model, preventing mask neglect or misinterpretation as noise.
@@
  • Proposing MedSegLatDiff for medical image segmentation, which simulates the segmentation process of a group of doctors by generating multiple outputs per input image to approximate their variance. Confidence maps derived from these outputs assist experts in performing more thorough and reliable diagnoses.
@@ II Related work
Medical Image Segmentation is the task of assigning a label to each pixel to determine its class membership. This problem has been extensively studied using various AI architectures. Early models following a one-to-one approach include fully convolutional networks [1] and encoder–decoder architectures with skip connections, such as UNet [2] and UNet++ [3], which extend skip connections for improved feature propagation. Traditional CNN-based models, including nnUNet [23] and ResUNet [31], have further demonstrated strong performance in medical image segmentation. In addition, transformer-based architectures, such as SegFormer [20], TransUNet [29], SwinUNet [30], have been applied for more comprehensive image analysis, and hybrid approaches combining hypernetworks, such as HyperSeg [28], have also been explored.
@@
Deep Generative Models are models initially developed for image synthesis tasks. Among them, VAE [14], which rely on probabilistic latent variable modeling, and GAN [19], which learn through adversarial training, laid the foundation. Meanwhile, DM [5], which produce images by gradually adding noise through a Markov chain process, have recently achieved state-of-the-art performance across various tasks, often matching or surpassing the most advanced GAN techniques [6, 4]. By incorporating conditioning into the generation process, several diffusion-based models have been applied to medical image segmentation. Specifically, EnsembleDiff [7] combines medical images with noisy images using a concatenation operation during denoising, while SegDiff [8] adds (rather than concatenates) the input image after it passes through a convolutional encoder. MedSegDiff [9] and MedSegDiff-v2 [10] further leverage the Feature Frequency Parser to mitigate high-frequency noise and employ the Spectrum-Space Transformer for semantic feature interaction. Furthermore, diffusion models have been integrated with VAEs, as in Stable Diffusion [11], performing forward and reverse processes in latent space to train the denoising model and synthesize new data. This approach has achieved success in image inpainting, class-conditional image synthesis, and other tasks, including text-to-image synthesis, unconditional image generation, and super-resolution [12, 13, 11]. Inspired by this, we adopt a latent diffusion approach for medical image segmentation, called MedSegLatDiff. Its stochastic sampling allows generation of multiple segmentation outputs per image, capturing variability similar to that of a group of doctors. This makes it well-suited for handling ambiguous and challenging medical data.
@@ III Method
Figure 1: The latent diffusion forward process, where noise is progressively added across diffusion steps.
The proposed model, MedSegLatDiff, integrates a conditional diffusion-based model with VAEs. This section first provides a brief overview of conditional diffusion models for medical image segmentation and VAEs model for encoding and decoding both medical images and segmentation masks. It then focuses on latent diffusion models (LDMs) with conditioning mechanisms operating in latent space for this task.
III-A Segmentation in image pace
Formally, we consider a medical dataset:
@@ III-A Segmentation in image pace
where $X^{(i)}\in\mathbb{R}^{H\times W\times C}$ denotes the $i$-th medical image, $S^{(i)}\in\{0,1\}^{H\times W}$ is its corresponding segmentation mask, and $N$ is the total number of image–mask pairs in the dataset. The main idea when applying DM for segmentation task is to gradually add noise $\epsilon\sim\mathcal{N}(0,I)$ to a mask $S$ through a forward process and progressively remove noise from the noisy image via a reverse process. The forward process starts with the original image $S_{0}$ sampled from the mask distribution $p_{\textrm{mask}}(S)$ and is defined as:
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$. This forward process generates noisy image pairs that serve as training data for a denoising model $\epsilon_{\theta}(S_{t},t)$. The model takes the noisy image $S_{t}$ as input and the added noise $\boldsymbol{\epsilon}$ at step $t$ as the target. It is trained using a simple Mean Squared Error (MSE) loss:
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}\left\|\epsilon-\epsilon_{\theta}(S_{t},t)\right\|_{2}^{2},$ (4)
and is based on a UNet architecture.
For the inference process, we approxiate the reveser distribution $q(S_{t-1}\mid S_{t},S_{0})$ by trainned denoising model. Therefore, starting from pure Gaussian noise $S_{T}\sim\mathcal{N}(0,I)$, the noise is iteratively reduced at each step to recover $S_{t-1}$. This is called as thes reverse process and performed as follows:
@@
where $\epsilon^{\prime}\sim\mathcal{N}(0,I)$ introduces stochasticity into the reverse process, and $t=T,\dots,1$ denotes the reverse diffusion steps. Therefore, by repeating the $T$ reverse steps, the model can generate multiple segmentation masks $S_{0}$, which approximate samples drawn from the original mask distribution.
However, the setup described above only generates arbitrary masks and cannot perform segmentation on an input image. Specifically, given a medical image XXand its corresponding segmentation mask $S$, we adapt the model for the segmentation task by appending $X$ as an additional input channel to incorporate anatomical information. Therefore, the mask and the input image are concatenated channel-wise to obtain
$S^{\mathrm{cond}}:=S\oplus X,$ (6)
where $\oplus$ denotes channel-wise concatenation. The resulting conditioned representation is then fed into the modified denoising model $\epsilon(S^{\mathrm{cond}},t)$. This concatenation guides the denoising process, ensuring that the model generates new masks corresponding to the input image $X$. During the forward process, noise is added only to the ground truth segmentation $M$, while $X$ remains unchanged to preserve anatomical guidance. Since the sampling process is stochastic, the diffusion model generates multiple segmentation masks $S_{0}$ for each input image $X$. When trained with a single segmentation mask per image, the model implicitly produces an ensemble of masks, effectively emulating the consensus-driven segmentation process of a group of doctors and potentially improving overall performance.
@@ III-B Image and Mask in latent space
The VAE [14] is a neural network that maps each image from the data distribution to a latent code in a lower-dimensional latent space. It consists of an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$. The encoder maps an image $X\in\mathbb{R}^{H\times W\times C}$ to a latent representation $z=\mathcal{E}(X)$, while the decoder reconstructs the image as $X^{\prime}=\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(X))$, where $z\in\mathbb{R}^{h\times w\times c}$ lies in a lower-dimensional latent space. The encoder downsamples the image by a factor $f=\frac{H}{h}=\frac{W}{w}$, and in our study, we investigate different downsampling factors of the form $f=2^{m}$ with $m\in\mathbb{N}$. For our segmentation task, both the medical image $X$ and the corresponding segmentation mask $S$ must be transformed into a compact latent space. To achieve this, we employ two separate autoencoders trained independently for images and masks. In particular, we adopt the Vector Quantized Variational Autoencoder (VQ-VAE) [24, 25], whose discrete latent space has demonstrated strong generative performance in recent models such as Stable Diffusion [11]. VQ-VAE is particularly suitable for our setting because: (i) its discrete codebook provides a stable, expressive latent representation; (ii) the quantized latent space is naturally compatible with diffusion models, as demonstrated in Stable Diffusion; and (iii) the stop-gradient mechanism stabilizes training and preserves the piecewise-constant structure characteristic of segmentation masks.
@@
Secondly, we design a separate VQ-VAE for the segmentation mask to enable the diffusion-based model to learn the mask distribution in the latent space more efficiently and effectively. It is structured similarly to the image VQ-VAE and consists of an encoder $\mathcal{E}_{S}(\cdot)$, a decoder $\mathcal{D}_{S}(\cdot)$, and a quantization module $Q_{\mathcal{V}_{S}}(\cdot)$, which discretizes the latent representation using a codebook $\mathcal{V}_{S}=\{v_{S}^{(k)}\}_{k=1}^{K}$. The key difference lies in our modified loss function, which focuses on the segmented region, with particular emphasis on tiny nodules in the LIDC-IDRI dataset. Instead of using MSE for reconstruction, we adopt WCE to better capture the characteristics of small and sparse lesion regions. The modified reconstruction loss assigns larger weights to pixels within the segmentation mask, enabling the model to focus on these regions and reconstruct the mask more accurately, while all other loss components remain unchanged. This design prevents the perceptual compression model from ignoring tiny nodules, which might otherwise be mistakenly treated as noise.
III-C Segmentation in latent spcce
Once prepared two pre-trained VQ-VAE models: one for medical images and the other for the corresponding segmentation masks follow by Part III-B, we perform the diffusion processes in the latent domain. For each pair of medical image $X\in\mathbb{R}^{H\times W\times C}$ and segmentation mask $S\in\mathbb{R}^{H\times W}$, the latent data is obtained $\bar{z}_{X}$ and $\bar{z}_{S}$ corresponding. Using the latent representations, the diffusion model performs the forward diffusion process on each latent mask $z_{S,0}=\bar{z}_{S}$, as illustrated in Fig. 1. This process follows the diffusion mechanism described in Eqs. 2 and 3, which generate noisy labels for the data and train the denoising model.

### v3

@@ Abstract
Medical image segmentation is essential for supporting clinical diagnosis and treatment planning. With growing imaging demands, artificial intelligence (AI) models are increasingly used to improve the accuracy and efficiency of this task while reducing the workload of radiologists. However, traditional approaches often produce a single segmentation mask per input image, limiting the ability to capture uncertainty in the segmentation process. Recent advancements in generative models for image generation present new opportunities to address these limitations by enabling the generation of multiple segmentation masks for each input image. Nevertheless, computational complexity and performance remain key challenges when operating directly in the image domain. In this work, we propose MedSegLatDiff, an effective framework for medical image segmentation that integrates a diffusion-based model (DM) with two variational autoencoders (VAEs) to decouple the segmentation process from perceptual data compression. Specifically, VAEs compress data into a low-dimensional latent space, reducing noise and accelerating learning, while the DM performs segmentation more efficiently within this latent representation. Furthermore, we replace the traditional Mean Squared Error (MSE) loss with a weighted Cross-Entropy (WCE) loss in the perceptual mask compression module, which enhances the reconstruction of segmentation masks during the encode–decode process, particularly for tiny nodules. Our designed framework emulates the collaborative segmentation approach of multiple clinicians, resulting in more robust and reliable performance compared to previous methods that model the segmentation patterns of individual practitioners. We evaluate our model on three datasets: ISIC-2018, CVC-Clinic, and LIDC-IDRI, demonstrating competitive performance while simultaneously providing the ability to generate confidence maps for deeper analysis by experts. Our results show that MedSegLatDiff achieves prominent segmentation performance while offering enhanced interpretability and consistency for clinical applications.
Keywords:
diffusion model; variational autoencoder; medical image segmentation; conditional image generation.
@@ I Introduction
Medical image segmentation, which involves partitioning medical images into meaningful regions, plays a crucial role in clinical applications such as diagnosis, surgical planning, and image-guided procedures [3]. Accurate segmentation enables clinicians to better interpret images, facilitate comparisons, and monitor changes over time. However, manually labeling segmentation masks is time-consuming and labor-intensive, motivating the development of automated methods. While classical techniques such as binary thresholding [6] and k-means clustering [4] have shown limited effectiveness [7, 5, 1], recent advances in AI have enabled the widespread adoption of machine learning (ML) and deep learning (DL) models, which achieve remarkable results [2].
@@
Typically, these AI models are trained on pre-labeled datasets and can automatically segment new images. Most existing approaches adopt a one-to-one strategy, producing only a single segmented output per input image [9, 10, 11, 12]. These methods basically learn a mapping function from each input image to a single segmentation mask, thereby replicating the behavior of an individual annotator. Recently, generative models, particularly diffusion models (DMs) [22], have led to substantial advances in image synthesis, often outperforming earlier approaches such as Generative Adversarial Networks (GANs) [21] and Variational Autoencoders (VAEs) [18], which remain among the most widely used methods in recent years. Beyond their success in natural image generation, these models open up new possibilities for medical image segmentation through a one-to-many paradigm. Unlike the previous one-to-one approach, diffusion-based models can generate multiple segmentation masks for a single input, capturing the model’s uncertainty in ambiguous medical data. Several recent studies have adopted this strategy and demonstrated superior performance and robustness compared to conventional methods [30, 31, 32, 38]. These results indicate that diffusion-based approaches not only enhance segmentation accuracy but also provide richer information to support clinical decision-making, particularly in challenging cases involving ambiguous or subtle anatomical structures. However, these models typically operate directly in the image space, where performance is limited because they are required to jointly perform image compression and region-of-interest analysis. Consequently, we identify opportunities to further improve model performance by conducting segmentation in a low-dimensional latent space.
@@
In this paper, we propose MedSegLatDiff, a medical image segmentation framework that adopts a novel one-to-many paradigm based on DM. Our framework integrates two VAEs with a diffusion-based model and consists of two main stages. The first stage performs an encoder–decoder architecture to map the input image into a low-dimensional latent representation, while the second stage performs diffusion processes in the latent space to generate a diverse set of segmentation masks for each input image. By conditioning the stochastic generation process of the diffusion model in the latent space, the model produces a set of segmentation masks corresponding to the input, whose variability captures the uncertainty typically exhibited by a group of clinicians. Thus, our model effectively simulates the segmentation consensus among multiple doctors and is well-suited for medical image segmentation tasks. Furthermore, the segmented results from the generated set of masks compete with those of previous models, while confidence maps derived from these outputs assist clinicians in making more thorough and reliable diagnoses. To the best of our knowledge, the main contributions of this paper are as follows:
@@
  • Training and incorporating VAEs to develop a conditional diffusion-based model in latent space, which reduces noise, compresses image into latent space, and improves both training and inference efficiency.
  • Replacing Mean Squared Error (MSE) with weighted Cross-Entropy (WCE) in the VAE mask reconstruction stage improves the preservation of small structures, especially in cases involving tiny masks, thereby reducing the risk of neglect or misinterpretation as noise.
@@
  • Proposing MedSegLatDiff for medical image segmentation, which simulates the segmentation process of a group of clinicians by generating multiple outputs per input image to approximate their variance. Futhermore, confidence maps derived from these outputs assist experts in performing more thorough and reliable diagnoses.
@@ II Related Work
Medical Image Segmentation refers to the process of assigning a label to each pixel (or voxel) in a medical image to determine its class membership, thereby identifying regions of interest with diverse shapes. This problem has been extensively studied using various AI architectures, demonstrating superior performance compared to classical methods [2]. Early models followed a one-to-one paradigm, evolving from fully convolutional networks (FCNs) [8], which employ conventional CNN-based architectures, to encoder–decoder architectures with skip connections, such as U-Net [9] and U-Net++ [10], which extend skip connections to improve feature propagation. Traditional CNN-based models, including nnU-Net [11] and ResUNet [12], have demonstrated strong performance in medical image segmentation by effectively capturing hierarchical features and leveraging multi-scale contextual information. However, these methods are limited in modeling long-range dependencies, which are essential for capturing global contextual information in complex medical images. To address this limitation, transformer-based architectures such as SegFormer [16], TransUNet [13], and SwinUNet [14] have been introduced to enable more comprehensive image analysis through deep transformer blocks at the cost of increased computational complexity and memory consumption. In addition, hybrid approaches that incorporate hypernetworks, such as HyperSeg [17], have been explored but remain complex and insufficiently mature. Overall, although these methods are still actively developing and show promising performance, they largely follow a one-to-one paradigm and are unable to explicitly model uncertainty in the segmentation process, which is crucial in medical applications.
@@
Deep Generative Models were initially developed for image generation and synthesis, but their ability to learn underlying data distributions has recently drawn increasing attention in medical image segmentation and uncertainty modeling. Among them, Variational Autoencoders (VAEs) [18], which rely on probabilistic latent variable modeling, and Generative Adversarial Networks (GANs) [21], which learn through adversarial training, laid the foundation for modern deep generative modeling. More recently, diffusion models (DMs) [22], which generate images by gradually adding and then removing noise through a Markov chain process, have achieved state-of-the-art performance across various tasks, often matching or surpassing advanced GAN techniques [25]. By incorporating conditioning into the generation process, several diffusion-based models have been successfully applied to various domains, including medical image segmentation. For example, EnsembleDiff [30] conditions the diffusion process by concatenating medical images with their noisy counterparts during denoising, whereas SegDiff [31] adds (rather than concatenates) the input image via element-wise addition after passing it through a convolutional encoder. Notably, MedSegDiff [32] and MedSegDiff-v2 [33] further introduce a Feature Frequency Parser to mitigate high-frequency noise and employ a Spectrum-Space Transformer to enhance semantic feature interaction. Other conditioning approaches include guided diffusion for universal anatomical segmentation [36] and prompt-guided variants such as ProGiDiff [37], which adapt pre-trained latent diffusion models for flexible conditioning. In parallel, applications of diffusion models have continued to evolve to alleviate the computational burden. Two major directions have emerged: performing the diffusion process in a low-dimensional latent space, as exemplified by Stable Diffusion [27], and accelerating the inference process through improved sampling strategies such as DDIM [24] and PLMS [43]. The former has proven effective for a wide range of tasks, including image inpainting, class-conditional synthesis, text-to-image generation, unconditional image generation, and super-resolution [28, 29, 27]. The latter direction has been widely explored in medical image segmentation. For instance, LSegDiff [34] proposes a fast latent diffusion model evaluated on polyp and skin lesion segmentation tasks, while LDSeg [35] introduces an end-to-end conditional latent diffusion framework that leverages low-dimensional shape manifolds for multi-modal segmentation, achieving faster sampling and reduced memory consumption than pixel-space diffusion methods.
While these prior works demonstrate the promise of latent diffusion in improving efficiency and accuracy for medical image segmentation, most rely on standard VAEs or pixel-space conditioning and do not explicitly address challenges such as tiny or sparse target structures (e.g., small nodules) through specialized loss designs, such as weighted reconstruction. Motivated by these observations, we propose MedSegLatDiff, a latent diffusion framework that integrates VQ-VAE [19, 20] for discrete and stable compression of both images and segmentation masks. In particular, a WCE reconstruction loss is incorporated into the mask VQ-VAE to replace the conventional MSE loss, thereby improving the preservation of small and sparse structures. Furthermore, stochastic sampling in the latent space enables a one-to-many paradigm that generates multiple plausible segmentation masks for a given input, capturing variability comparable to that observed among groups of clinicians. As a result, MedSegLatDiff is particularly well suited for ambiguous and challenging medical data, where uncertainty modeling can provide valuable support for clinical decision-making.
@@ III Method
Figure 1: Illustration of the forward diffusion process in latent space, in which Gaussian noise is progressively added to the encoded segmentation mask across diffusion timesteps.
In this section, we introduce our proposed MedSegLatDiff framework, which integrates a conditional diffusion-based model with two VAEs for medical image segmentation. We begin with an overview of conditional diffusion models for image-space segmentation, followed by a description of the VAEs used to encode and decode medical images and segmentation masks. We then present our latent diffusion model with conditioning mechanisms operating in the latent space for this task.
III-A Segmentation in image space
Medical image segmentation is formulated under the assumption that a labeled dataset is available, where each image is paired with a ground-truth segmentation mask. Formally, we consider a medical imaging dataset:
@@ III-A Segmentation in image space
where $X^{(i)}\in\mathbb{R}^{H\times W\times C}$ denotes the $i$-th medical image, $S^{(i)}\in\{0,1\}^{H\times W}$ is its corresponding segmentation mask, and $N$ is the total number of image–mask pairs in this dataset $\mathcal{D}$. The main idea when applying DM for segmentation task is to gradually add noise $\epsilon\sim\mathcal{N}(0,I)$ to a mask $S$ through a forward process and progressively remove noise from the noisy image via a reverse process. The forward process starts with the original image $S_{0}$ sampled from the mask distribution $p_{\textrm{mask}}(S)$ and is defined as:
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$. This forward process generates noisy image pairs that serve as training data for a denoising model $\epsilon_{\theta}(S_{t},t)$. The model takes the noisy image $S_{t}$ as input and the added noise $\boldsymbol{\epsilon}$ at step $t$ as the target. In its simplest form, the model is trained using the MSE loss:
$\mathcal{L}_{\textrm{simple}}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}\left\|\epsilon-\epsilon_{\theta}(S_{t},t)\right\|_{2}^{2},$ (4)
and the noise estimator $\epsilon_{\theta}(S_{t},t)$ is implemented using a U-Net architecture [22].
During inference, the reverse distribution $q(S_{t-1}\mid S_{t},S_{0})$ is approximated using the trained denoising model. Starting from pure Gaussian noise $S_{T}\sim\mathcal{N}(0,I)$, the noise is iteratively removed at each timestep to recover $S_{t-1}$. This procedure is referred to as the reverse diffusion process and is performed as follows:
@@
where $\epsilon^{\prime}\sim\mathcal{N}(0,I)$ introduces stochasticity into the reverse process, and $t=T,\dots,1$ denotes the reverse diffusion steps. By repeating the reverse process for $T$ timesteps, the model generates a segmentation mask $S_{0}$ as an approximate sample from the underlying distribution of segmentation masks. Moreover, repeated stochastic sampling enables the generation of multiple plausible segmentation masks, effectively approximating this distribution.
However, the formulation described above only generates unconditional random masks and cannot perform segmentation conditioned on a given input image. To address this limitation, following prior studies, we incorporate conditioning into the generation process, enabling the model to generate a corresponding segmentation mask for a given input image. Specifically, given a medical image $X$ and its corresponding segmentation mask $S$, we adapt the model for the segmentation task by appending $X$ as an additional input channel to incorporate anatomical information into each noisy mask $S_{t}$. As a result, the noisy mask and the input image are concatenated channel-wise to obtain
$S_{t}^{\mathrm{cond}}:=S_{t}\oplus X,$ (6)
where $\oplus$ denotes channel-wise concatenation. The resulting conditioned representation is then fed into the modified denoising model $\epsilon(S_{t}^{\mathrm{cond}},t)$. This concatenation guides the denoising process, ensuring that the model generates new masks corresponding to the input image $X$. During the forward process, noise is added only to the ground truth segmentation $S$, while $X$ remains unchanged to preserve anatomical guidance. Since the sampling process is stochastic, the diffusion model generates multiple segmentation masks $S_{0}$ for each input image $X$. When trained with a single segmentation mask per image, the model implicitly produces an ensemble of masks, effectively emulating the consensus-driven segmentation process of a group of doctors and potentially improving overall performance.
@@ III-B Image and Mask in latent space
The VAE [18] is a neural network that maps each image from the data distribution to a latent code in a lower-dimensional latent space. It consists of an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$. The encoder maps an image $X\in\mathbb{R}^{H\times W\times C}$ to a latent representation $z=\mathcal{E}(X)$, while the decoder reconstructs the image as $X^{\prime}=\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(X))$, where $z\in\mathbb{R}^{h\times w\times c}$ lies in a lower-dimensional latent space. The encoder downsamples the image by a factor $f=\frac{H}{h}=\frac{W}{w}$, and in our study, we investigate different downsampling factors of the form $f=2^{m}$ with $m\in\mathbb{N}$. For our segmentation task, both the medical image $X$ and the corresponding segmentation mask $S$ must be transformed into the same low-dimensional latent space. To achieve this, we employ two separate autoencoders trained independently for images and masks. In particular, we adopt the Vector Quantized Variational Autoencoder (VQ-VAE) [19, 20], whose discrete latent space has demonstrated strong generative performance in recent models such as Stable Diffusion [27]. VQ-VAE is particularly suitable for our setting because: (i) its discrete codebook provides a stable, expressive latent representation; (ii) the quantized latent space is naturally compatible with diffusion models, as demonstrated in Stable Diffusion [27]; and (iii) the stop-gradient mechanism stabilizes training and preserves the piecewise-constant structure characteristic of segmentation masks.
@@
Secondly, we design another VQ-VAE for the segmentation mask to enable the diffusion-based model to learn the mask distribution in the latent space more efficiently and effectively. It is structured similarly to the image VQ-VAE and consists of an encoder $\mathcal{E}_{S}(\cdot)$, a decoder $\mathcal{D}_{S}(\cdot)$, and a quantization module $Q_{\mathcal{V}_{S}}(\cdot)$, which discretizes the latent representation using a codebook $\mathcal{V}_{S}=\{v_{S}^{(k)}\}_{k=1}^{K}$. The key difference lies in our modified loss function, which focuses on the segmented region, with particular emphasis on tiny nodules in the LIDC-IDRI dataset. Instead of using MSE for reconstruction, we adopt WCE to better capture the characteristics of small and sparse lesion regions. The modified reconstruction loss assigns larger weights to pixels within the segmentation mask, enabling the model to focus on these regions and reconstruct the mask more accurately, while all other loss components remain unchanged. This design prevents the perceptual compression model from ignoring tiny nodules, which might otherwise be mistakenly treated as noise.
III-C Segmentation in latent space
Figure 2: Overview of segmentation pipeline: multiple segmentation masks are generated from random noise in the latent space for a given medical image, followed by decoding, consensus fusion, and confidence map estimation.
Once the two pre-trained VQ-VAE models, one for medical images and the other for the corresponding segmentation masks, are prepared as described in Part III-B, we perform the diffusion processes in the latent domain. For each pair of medical image $X\in\mathbb{R}^{H\times W\times C}$ and segmentation mask $S\in\mathbb{R}^{H\times W}$, the corresponding latent representations are obtained as $\bar{z}_{X}$ and $\bar{z}_{S}$. Using the latent representations, the diffusion model performs the forward diffusion process on each latent mask $z_{S,0}=\bar{z}_{S}$, as illustrated in Fig. 1. This process follows the diffusion mechanism described in Eqs. 2 and 3, which generate noisy labels for the data and train the denoising model.